# Properties of observable mixed inertial/gravito-inertial modes in $\gamma$ Doradus stars

### v1

@@ 2.4 Numerical resolution
In this subsection, we estimate which radial and latitudinal resolutions are needed for our analysis. The radial resolution is determined by the number of radial points $n_{r}$ and needs to be adjusted depending on the number of radial nodes of the considered mode.
The latitudinal resolution $n_{\theta}$ corresponds to the number of associated Legendre polynomials $P_{\ell}^{m}$ on which the modes are decomposed. The degree $\ell$ span from $\ell_{\mathrm{min}}=|m|+i_{p}$ to $\ell_{\mathrm{max}}=\ell_{\rm min}+2(n_{\theta}-1)$, where $i_{p}=0$ or 1 when $\ell+m$ is even or odd respectively. For example, a mode of the series ($\ell=1$, $m=-1$) calculated with $n_{\theta}=3$ is decomposed latitudinally on the three first associated Legendre polynomials with the same azimuthal number and equatorial parity $P_{1}^{-1}$, $P_{3}^{-1}$ and $P_{5}^{-1}$. The minimal latitudinal resolution needed to compute a given mode can be estimated by looking for the number of associated Legendre polynomials needed to represent the associated Hough function. Figure 2 shows that for the Hough function $H_{\ell=1}^{m=-1}$ calculated at the spin parameter $s=2\Omega/\omega_{\mathrm{co}}\sim 16.4$ the projection on Legendre polynomials peaks at $P_{1}^{-1}$ while the contributions of polynomials above $P_{11}^{-1}$ are very small.This suggests that a latitudinal resolution of at least $n_{\theta}=6$ is needed for this mode.
@@
For a more accurate estimation, we apply the same process with modes calculated by TOP. For example, in the upper panel of Fig. 3, we see the decomposition of the mean latitudinal profile of a mode of the series ($\ell=1,m=-1$) onto associated Legendre polynomials, for different latitudinal resolutions. We first notice that the amplitude decreases rapidly with the degree $\ell_{i}$. Nevertheless, when we use too few polynomials, we miss non-negligible contributions of polynomials of degree $\ell_{i}$ larger than $\ell_{\mathrm{max}}$, leading to inaccuracies. It translates into errors on the frequency determination. In the lower panel of the same figure, we see how the computed frequency evolves with the latitudinal resolution and converges toward a stable value at large $\ell_{\mathrm{max}}$. For this mode, the frequency varies by less than 0.01% for $\ell_{\mathrm{max}}$ higher than 9 ($n_{\theta}=5$).
@@
Figure 3: Upper panel: projections of the mean latitudinal profile in the radiative zone of the mode $n=66$ ($s\sim 16.4$), from the series of modes ($\ell=1$, $m=-1$) at different latitudinal resolution $n_{\theta}$ on associated Legendre polynomials for a rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$ with $\Omega_{\mathrm{K}}=\sqrt{GM/R^{3}}$ the critical rotation rate of the star, $M$ and $R$ being its mass and radius respectively. Lower panel: evolution of the relative frequency of the same mode as a function of $\ell_{\mathrm{max}}$. The reference frequency $\omega_{\mathrm{ref}}$ corresponds to the frequency obtained using the highest latitudinal resolution ($n_{\theta}=12$). The model used is 1z.
@@
As we study dips, the resolution will be chosen such that the dips are computed with enough accuracy. Figures 4 and 5 show the impact of the radial and latitudinal resolutions on a series of modes, in diagrams depicting period spacing $\Delta P_{\mathrm{co}}$ (period differences of two modes with consecutive $n$ in the co-rotating frame) as a function of the spin parameter $s$ for the rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. The vertical line indicates the spin parameter of the expected resonant mode in the case of a core of uniform density (see Sect. 3.2). We see that the dip deepens and that $\Delta P_{\mathrm{co}}$ increases for the higher $s$ as the radial resolution decreases, which increases the general slope of the curve. When the latitudinal resolution decreases, $\Delta P_{\mathrm{co}}$ decreases for the higher $s$, which leads to the opposite change in the general slope of the curve. The curve barely evolves for $n_{r}>1000$ and $n_{\theta}>6$, which confirms the primary guess we made using the projection of the Hough function on the associated Legendre polynomials. Based on Fig. 4 and knowing that the mode with the highest number of radial nodes is $n=72$, we deduce that about $15\times n$ points is enough to reach convergence.
@@
Figure 4: Period spacing $\Delta P_{\mathrm{co}}$, normalised by $\Pi_{0}$ (buoyancy radius), plotted as a function of the spin parameter $s$ with different radial resolution $n_{r}$, for the series of modes ($\ell=1$, $m=-1$) and for a rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. Their radial order $n$ goes from $32$ to $72$. The model used is 1z. The pink line shows the traditional approximation of rotation. The vertical line shows the spin parameter of the resonant inertial mode calculated with an analytical uniform-density model.
@@
Figure 5: Period spacing $\Delta P_{\mathrm{co}}/\Pi_{0}$ plotted as a function of the spin parameter $s$ with different latitudinal resolutions $n_{\theta}$, for the series of modes ($\ell=1$, $m=-1$) and for a rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. Their radial order $n$ goes from $32$ to $72$. The model used is 1z. The pink line shows the traditional approximation of rotation. The vertical line shows the spin parameter of the resonant inertial mode calculated with an analytical uniform-density model.
@@
To describe the dips accurately, we want the error on the period spacing $\delta\Delta P_{\mathrm{co}}$ not to exceed 1% of the total depth of the dip. We determined the radial and latitudinal resolutions for our numerical calculations to meet on this criterion on the frequency error $\delta\omega$ using
@@
To ensure that $\delta\omega<1\%$, we showed it is sufficient to chose a radial resolution $n_{r}$ ranging from 5000 to 10000 for the ($\ell=1$, $m=-1$) series, from 3000 to 10000 for the ($\ell=2$, $m=-2$) series and from 3000 to 10000 for the ($\ell=1$, $m=0$) series. These ranges cover the different resolutions that are needed, since the radial order increases with the spin parameter and evolutionary status. For the latitudinal resolution, we typically use $n_{\theta}=6$ to $8$ for the series ($\ell=1$, $m=-1$) and ($\ell=2$, $m=-2$), and 4 to 6 for the series ($\ell=1$, $m=0$).
Figure 6: (Left panel) Red dots represent the period spacing $\Delta P_{\mathrm{co}}$ and the pink line shows the TAR, both as a function of the spin parameter $s$ for the series ($\ell=1$, $m=-1$) for the 1z model at a rotation rate of $\Omega=0.1\Omega_{\mathrm{K}}$. The blue line shows the proportion of kinetic energy in the core for each mode. (Middle panel) Quantity $p/\rho_{0}^{0.8}$ plotted in a meridional plane for the mode marked by a black star in the left panel. (Right panel) Radial profile of $p/\rho_{0}^{0.8}$ at the latitude $\pi$/2 (equator). The factor $\rho_{0}^{0.8}$ is a scaling factor used to make the visualisation easier. The red vertical lines shows the radiative zone limits.
@@ 3 Results
In this section, we present the results of our numerical calculations. We studied the series of modes ($\ell=1$, $m=-1$), ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) at rotation rates ranging from $0.1$ to $0.5$ $\Omega/\Omega_{\mathrm{K}}$, with $\Omega_{\mathrm{K}}=\sqrt{GM/R^{3}}$, for the three stellar models described in Table 1. Among these three series, the ($\ell=1$, $m=-1$) is by far the most frequently observed (Li et al., 2020). For each series, we compute the period spacing $\Delta P_{\mathrm{co}}$ in a frequency interval where the gravito-inertial modes strongly interact with an inertial mode of the convective core. We then describe the properties of the dips obtained in the $\Delta P_{\mathrm{co}}$-$s$ diagram. In the following, we first present the basic characteristics of a typical dip in Sect. 3.1. The simple model we use to determine the approximate spin parameter of the dips is described in in Sect. 3.2. The results of our investigation of dips of the ($\ell=1$, $m=-1$), ($\ell=2$, $m=-2$) and ($\ell=1$, $m=0$) mode series are then reported in Sect. 3.3.
3.1 Basic dips properties
@@
We first describe the properties of a typical dip for the series ($\ell=1$, $m=-1$). The dip is located around $s=10.7$ and is calculated with the 1z model at rotation rate $\Omega=0.1\Omega_{\mathrm{K}}$ (shown in the left panel of Fig. 6). We can see the difference between our calculations and the TAR shown in pink: while there is a clear dip in the $\Delta P_{\mathrm{co}}-s$ curve around $s=10.7$ in our calculations, the TAR is nearly constant. For prograde sectoral modes ($\ell=-m$), the TAR indeed leads to a constant $\Delta P_{\mathrm{co}}$ in the limit of high spin parameters. Middle and right panels of Fig. 6 show the Eulerian pressure perturbation $p$ for the mode located at the centre of the dip. It exhibits a very high amplitude in the convective core compared to the radiative zone. The presence of significant oscillations in both the convective core and the radiative zone indicates the mixed nature of the mode. To confirm the mixed nature of the modes in the dip, we calculated the portion of kinetic energy in the core for each mode. It reads:
$\frac{E_{\mathrm{kin}}(\mathrm{core})}{E_{\mathrm{kin}}(\mathrm{core+RZ})}=\frac{\int_{0}^{r_{c}}\int_{0}^{2\pi}\int_{0}^{\pi}\rho_{0}\left|\@vec{v}\right|^{2}r^{2}\sin{\theta}\,\mathrm{d}\theta\,\mathrm{d}\phi\,\mathrm{d}r}{\int_{0}^{R}\int_{0}^{2\pi}\int_{0}^{\pi}\rho_{0}\left|\@vec{v}\right|^{2}r^{2}\sin{\theta}\,\mathrm{d}\theta\,\mathrm{d}\phi\,\mathrm{d}r},$ (8)
@@
$E_{\mathrm{kin}}(\mathrm{core})$ and $E_{\mathrm{kin}}(\mathrm{core+RZ})$ being the mode kinetic energy in the core and the mode kinetic energy in the core and the radiative zone respectively. The mode perturbation velocity is denoted ${v}$, the radius of the core $r_{c}$ and the radius of the star $R$. As observed in the left panel of Fig. 6, the portion of kinetic energy in the core increases for the modes in the dip, reaching its maximum when $\Delta P_{\mathrm{co}}$ is minimum. Dips are formed by mixed gravito-inertial/inertial modes around the spin paremeter of an inertial mode of the core that couples with the gravito-inertial modes of the $(\ell,m)$ series. In the dip, a frequency is added to the the series.
@@
To study the dips and quantify their evolution, we model them as inverse Lorentzian profiles, following Tokuno & Takata (2022):
@@
where $s$ is the spin parameter of the gravito-inertial modes in the co-rotating frame, and $x_{1}$, $x_{2}$, $s_{c}$ and $\sigma$ are free fitting parameters. With this parametrisation, $x_{1}$ and $x_{2}$ account for the general slope of the curve, $s_{c}$ is the spin parameter on which the dip is centred and $\sigma$ is the width of the dip.
Adding one frequency locally modifies the period spacing but not the whole period gap between unaffected modes. This constraint imposes that the integral of the Lorenztian term over $s$ is constant and thus that the width and the depth are anti-correlated (Tokuno & Takata, 2022). We fit this model to our numerically computed dip using a least-square method. The errors associated with the fit tend to increase with the rotation rate, as the dip implies less modes when the rotation increases, decreasing the quality of the fit. The relative error of $s_{c}$ is negligible, the one of $\sigma$ is always inferior to 1% for the series ($\ell=1$, $m=-1$), to 2% series ($\ell=2$, $m=-2$) but goes up to 10% with the rotation rate for the series ($\ell=1$, $m=0$). The latter case could be explained by the fact that $x_{1}$ and $x_{2}$ account for a linear slope of $1/\Delta P_{\mathrm{co}}$, which is only a good local approximation of the TAR. As the rotation rate increases, the dip widens and this approximation becomes less valid, explaining the growing relative error.
@@
Assuming a convective core of uniform density, Ouazzani et al. (2020) proposed a simple analytical model to identify the inertial modes that produce significant dips in the period spacings of gravito-inertial mode series. Ouazzani et al. (2020) showed that this model provides a useful approximation of the dip spin parameters observed in full numerical computations.
We thus used this model to choose the dips we want to study and to get an initial guess of their spin parameter. The model details are given in Appendix A. Once a dip is found numerically, we can easily check that it is due to the expected inertial mode because the spatial structure of the inertial mode in our complete calculations (see for example the core region in the middle panel of Fig. 6) is very similar to that of the inertial mode in the sphere of uniform density (see comparisons in Appendix A).
For each series of gravito-inertial modes, we limited our study to resonances that could occur in spin-parameter ranges actually observed in $\gamma$ Dor stars. This condition is verified for one resonance in each series, which happens to be the one occurring at the smallest spin parameter.

### v2

@@ 2.4 Numerical resolution
In this subsection, we estimate which radial and latitudinal resolutions are needed for our analysis. The radial resolution was determined by the number of radial points $n_{r}$ and needed to be adjusted depending on the number of radial nodes of the considered mode.
The latitudinal resolution $n_{\theta}$ corresponds to the number of associated Legendre polynomials $P_{\ell}^{m}$ on which the modes are decomposed. The degree $\ell$ spans from $\ell_{\mathrm{min}}=|m|+i_{p}$ to $\ell_{\mathrm{max}}=\ell_{\rm min}+2(n_{\theta}-1)$, where $i_{p}=0$ or 1 when $\ell+m$ is even or odd, respectively. For example, a mode of the series ($\ell=1$, $m=-1$) calculated with $n_{\theta}=3$ is decomposed latitudinally on the three first associated Legendre polynomials with the same azimuthal number and equatorial parity $P_{1}^{-1}$, $P_{3}^{-1}$ and $P_{5}^{-1}$. The minimal latitudinal resolution needed to compute a given mode can be estimated by looking for the number of associated Legendre polynomials needed to represent the associated Hough function. Figure 2 shows that for the Hough function $H_{\ell=1}^{m=-1}$ calculated at the spin parameter $s=2\Omega/\omega_{\mathrm{co}}\sim 16.4$, the projection on Legendre polynomials peaks at $P_{1}^{-1}$, while the contributions of polynomials above $P_{11}^{-1}$ are very small.This suggests that a latitudinal resolution of at least $n_{\theta}=6$ is needed for this mode.
@@
For a more accurate estimation, we applied the same process with modes calculated by TOP. For example, in the upper panel of Fig. 3, one can see the decomposition of the mean latitudinal profile of a mode of the series ($\ell=1,m=-1$) onto the associated Legendre polynomials for different latitudinal resolutions. We first noticed that the amplitude decreases rapidly with the degree $\ell_{i}$. Nevertheless, when we used too few polynomials, we missed non-negligible contributions of polynomials of degree $\ell_{i}$ larger than $\ell_{\mathrm{max}}$, leading to inaccuracies. This translated into errors on the frequency determination. In the lower panel of the same figure, one can see how the computed frequency evolves with the latitudinal resolution and converges towards a stable value at a large $\ell_{\mathrm{max}}$. For this mode, the frequency varies by less than 0.01% for an $\ell_{\mathrm{max}}$ higher than nine ($n_{\theta}=5$).
@@
Figure 3: Effect of the latitudinal resolution on the mode $n=66$ ($s\sim 16.4$) of the series ($\ell=1$, $m=-1$). Upper panel: Projections of the mean latitudinal profile in the radiative zone of this mode at different latitudinal resolution $n_{\theta}$ on the associated Legendre polynomials for a rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. Here, $\Omega_{\mathrm{K}}=\sqrt{GM/R^{3}}$ is the critical rotation rate of the star and $M$ and $R$ are its mass and radius, respectively. Lower panel: Evolution of the relative frequency of the same mode as a function of $\ell_{\mathrm{max}}$. The reference frequency $\omega_{\mathrm{ref}}$ corresponds to the frequency obtained using the highest latitudinal resolution ($n_{\theta}=12$). The model used is 1z.
@@
Because we studied the dips, the resolution was chosen such that the dips are computed with enough accuracy. Figures 4 and 5 show the impact of the radial and latitudinal resolutions on a series of modes in diagrams depicting period spacing $\Delta P_{\mathrm{co}}$ (period differences of two modes with consecutive $n$ in the co-rotating frame) as a function of the spin parameter $s$ for the rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. The vertical line indicates the spin parameter of the expected resonant mode in the case of a core of uniform density (see Sect. 3.2). We observed that the dip deepens and that $\Delta P_{\mathrm{co}}$ increases for the higher $s$ as the radial resolution decreases, which increases the general slope of the curve. When the latitudinal resolution decreases, $\Delta P_{\mathrm{co}}$ decreases for the higher $s$, which leads to the opposite change in the general slope of the curve. The curve barely evolves for $n_{r}>1000$ and $n_{\theta}>6$, which confirms the primary guess we made using the projection of the Hough function on the associated Legendre polynomials. Based on Fig. 4 and knowing that the mode with the highest number of radial nodes is $n=72$, we deduced that about $15\times n$ points is enough to reach convergence.
@@
Figure 4: Period spacing $\Delta P_{\mathrm{co}}$, normalised by $\Pi_{0}$ (buoyancy radius), plotted as a function of the spin parameter $s$ with different radial resolution $n_{r}$ for the series of modes ($\ell=1$, $m=-1$) and for a rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. Their radial order $n$ goes from $32$ to $72$. The model used is 1z. The pink line shows the traditional approximation of rotation. The vertical line shows the spin parameter of the resonant inertial mode calculated with an analytical uniform density model.
@@
Figure 5: Period spacing $\Delta P_{\mathrm{co}}/\Pi_{0}$ plotted as a function of the spin parameter $s$ with different latitudinal resolutions $n_{\theta}$ for the series of modes ($\ell=1$, $m=-1$) and for a rotation rate $\Omega=0.4\Omega_{\mathrm{K}}$. Their radial order $n$ goes from $32$ to $72$. The model used is 1z. The pink line shows the traditional approximation of rotation. The vertical line shows the spin parameter of the resonant inertial mode calculated with an analytical uniform density model.
@@
To describe the dips accurately, we wanted the error on the period spacing $\delta\Delta P_{\mathrm{co}}$ not to exceed 1% of the total depth of the dip. We determined the radial and latitudinal resolutions for our numerical calculations in order to meet this criterion on the frequency error $\delta\omega$ using
@@
To ensure that $\delta\omega<1\%$, we showed that it is sufficient to choose a radial resolution $n_{r}$ ranging from 5000 to 10000 for the ($\ell=1$, $m=-1$) series, from 3000 to 10000 for the ($\ell=2$, $m=-2$) series, and from 3000 to 10000 for the ($\ell=1$, $m=0$) series. These ranges cover the different resolutions that are needed since the radial order increases with the spin parameter and evolutionary status. For the latitudinal resolution, we typically used $n_{\theta}=6$ to $8$ for the series ($\ell=1$, $m=-1$) and ($\ell=2$, $m=-2$), and $n_{\theta}=4$ to $6$ for the series ($\ell=1$, $m=0$).
Figure 6: Example of a resonant mode. Left panel: Period spacing $\Delta P_{\mathrm{co}}$ as a function of the spin parameter for the series ($\ell=1$, $m=-1$) and for the 1z model at a rotation rate of $\Omega=0.1\Omega_{\mathrm{K}}$. Red dots represent computed modes. The pink line shows the TAR. The blue line shows the proportion of kinetic energy in the core for each mode. Middle panel: Quantity $p/\rho_{0}^{0.8}$ plotted in a meridional plane for the mode marked by a black star in the left panel. Right panel: Radial profile of $p/\rho_{0}^{0.8}$ at the latitude $\pi$/2 (equator). The factor $\rho_{0}^{0.8}$ is a scaling factor used to make the visualisation easier. The red vertical lines show the radiative zone limits.
@@ 3 Results
In this section, we present the results of our numerical calculations. We studied the series of modes ($\ell=1$, $m=-1$), ($\ell=2$, $m=-2$), and ($\ell=1$, $m=0$) at rotation rates ranging from $0.1$ to $0.5$ $\Omega/\Omega_{\mathrm{K}}$, with $\Omega_{\mathrm{K}}=\sqrt{GM/R^{3}}$, for the three stellar models described in Table 1. Among these three series, the ($\ell=1$, $m=-1$) is by far the most frequently observed (Li et al., 2020). For each series, we computed the period spacing $\Delta P_{\mathrm{co}}$ in a frequency interval where the gravito-inertial modes strongly interact with an inertial mode of the convective core. We then described the properties of the dips obtained in the $\Delta P_{\mathrm{co}}-s$ diagram. In the following, we first present the basic characteristics of a typical dip in Sect. 3.1. The simple model we used to determine the approximate spin parameter of the dips is described in Sect. 3.2. The results of our investigation of dips of the ($\ell=1$, $m=-1$), ($\ell=2$, $m=-2$), and ($\ell=1$, $m=0$) mode series are then reported in Sect. 3.3.
3.1 Basic dip properties
@@
We first describe the properties of a typical dip for the series ($\ell=1$, $m=-1$). The dip is located around $s=10.7$ and was calculated with the 1z model at rotation rate $\Omega=0.1\Omega_{\mathrm{K}}$ (shown in the left panel of Fig. 6). One can see the difference between our calculations and the TAR shown in pink. While there is a clear dip in the $\Delta P_{\mathrm{co}}-s$ curve around $s=10.7$ in our calculations, the TAR is nearly constant. For prograde sectoral modes ($\ell=-m$), the TAR indeed leads to a constant $\Delta P_{\mathrm{co}}$ in the limit of high spin parameters. The middle and right panels of Fig. 6 show the Eulerian pressure perturbation $p$ for the mode located at the centre of the dip. It exhibits a very high amplitude in the convective core compared to the radiative zone. The presence of significant oscillations in both the convective core and the radiative zone indicates the mixed nature of the mode. To confirm the mixed nature of the modes in the dip, we calculated the portion of kinetic energy in the core for each mode. It reads
$\frac{E_{\mathrm{kin}}(\mathrm{core})}{E_{\mathrm{kin}}(\mathrm{core+RZ})}=\frac{\int_{0}^{r_{c}}\int_{0}^{2\pi}\int_{0}^{\pi}\rho_{0}\left|\vec{v}\right|^{2}r^{2}\sin{\theta}\,\mathrm{d}\theta\,\mathrm{d}\phi\,\mathrm{d}r}{\int_{0}^{R}\int_{0}^{2\pi}\int_{0}^{\pi}\rho_{0}\left|\vec{v}\right|^{2}r^{2}\sin{\theta}\,\mathrm{d}\theta\,\mathrm{d}\phi\,\mathrm{d}r},$ (8)
@@
with $E_{\mathrm{kin}}(\mathrm{core})$ being the mode kinetic energy in the core and $E_{\mathrm{kin}}(\mathrm{core+RZ})$ as the mode kinetic energy in the core and the radiative zone. The mode perturbation velocity is denoted as $\vec{v}$, the radius of the core as $r_{c}$, and the radius of the star as $R$. As observed in the left panel of Fig. 6, the portion of kinetic energy in the core increases for the modes in the dip, reaching its maximum when $\Delta P_{\mathrm{co}}$ is minimum. Dips are formed by mixed (inertial and gravito-inertial) modes around the spin parameter of an inertial mode of the core that couples with the gravito-inertial modes of the $(\ell,m)$ series. In the dip, a frequency is added to the series.
@@
To study the dips and quantify their evolution, we modelled them as inverse Lorentzian profiles by following Tokuno & Takata (2022):
@@
where $s$ is the spin parameter of the gravito-inertial modes in the co-rotating frame and $x_{1}$, $x_{2}$, $s_{c}$, and $\sigma$ are free fitting parameters. With this parametrisation, $x_{1}$ and $x_{2}$ account for the general slope of the curve, $s_{c}$ is the spin parameter on which the dip is centred, and $\sigma$ is the width of the dip.
Adding one frequency locally modifies the period spacing but not the whole period gap between unaffected modes. This constraint imposes that the integral of the Lorenztian term over $s$ is constant and thus that the width and the depth are anti-correlated (Tokuno & Takata, 2022). We fit this model to our numerically computed dip using a least-squares method. The errors associated with the fit tend to increase with the rotation rate, as the dip implies less modes when the rotation increases, decreasing the quality of the fit. The relative error of $s_{c}$ is negligible; the one of $\sigma$ is always less than 1% for the series ($\ell=1$, $m=-1$) and less than 2% for the series ($\ell=2$, $m=-2$), but it goes up to 10% with the rotation rate for the series ($\ell=1$, $m=0$). The latter case could be explained by the fact that $x_{1}$ and $x_{2}$ account for a linear slope of $1/\Delta P_{\mathrm{co}}$, which is only a good local approximation of the TAR. As the rotation rate increases, the dip widens, and this approximation becomes less valid, explaining the growing relative error.
@@
Assuming a convective core of uniform density, Ouazzani et al. (2020) proposed a simple analytical model to identify the inertial modes that produce significant dips in the period spacings of a gravito-inertial mode series. Ouazzani et al. (2020) showed that this model provides a useful approximation of the dip spin parameters observed in full numerical computations.
We thus used this model to choose the dips we wanted to study and to get an initial guess of their spin parameter. The model details are given in Appendix A. Once a dip was found numerically, we could easily check that it is due to the expected inertial mode because the spatial structure of the inertial mode in our complete calculations (see for example the core region in the middle panel of Fig. 6) is very similar to that of the inertial mode in the sphere of uniform density (see comparisons in Appendix A).
For each series of gravito-inertial modes, we limited our study to resonances that could occur in spin-parameter ranges actually observed in $\gamma$ Dor stars. This condition was verified for one resonance in each series, which happens to be the one occurring at the smallest spin parameter.